New Oscilliscope Part 2 – Repair

Continuing work on the TDS544A, I got all of the electrolytic capacitors off, and cleaned the board with simple green and alcohol.20130701_211313 I took careful notes about what value of capacitor went where, and how all of the cables are hooked up internally. I also noted any switch settings I could find.20130701_204417

The Digikey order of capacitors and a brand new ultra fine tip for my iron arrived.20130701_211357Here we are mid process on getting all the capacitors back on. It turns out I did not order enough of the 33µF capacitors, so I ended up substituting the 10µF capacitors in some places on the logic board, and on the front panel. The Acquisition board got all the proper values, since it might be important for the RAM and other major signal processing elements.20130701_220535

On the logic board, I replaced about half of the capacitors on the power rails with the smaller 10µF value.  Anything that didn’t appear to be on the main rail got the full 33µF when possible. I do have more 33s coming, but I couldn’t wait.

I reassembled everything except the outer case, and it passed all the power on tests!

With everything put back together, I borrowed the probes from my older TEK and started calibrating them. I also repaired a ground clip while I was at it, and also calibrated a third non-Tek probe that I had. It has no markings (except the 600VDC Cat1 safety markings), and the capacitor for calibration was in the probe tip, not in the BNC connector.
20130702_222934

With everything all tuned up, I ran a Signal Path Calibration from the utility menu, and let it do it’s thing for a while. I played a bit with some of the trigger settings, and started getting used to the menus. It will be nice to have a real scope for future projects. I may have to dig out my old USB floppy drive so I can get screenshots off it. Unfortunately, it uses a FFC to connect the floppy drive, so I don’t think I can pop in an USB floppy emulator easily.

When the other capacitors arrive Monday, I’ll consider changing some of the 10µF to 33µF, but I think everything will be fine now.

Here are all the failed caps removed:20130702_233030

Leave a Reply

Your email address will not be published. Required fields are marked *